# Multi-view Graph Structural Representation Learning via Graph Coarsening

### v1

@@ 4.2 Loop and Clique Coarsening
22:  return $G_{co}$
@@
Coarsening requires counting all loops and cliques in the graph. However, according to the definitions above, loops are contained within cliques. Moreover, since the Maximum Clique Problem (MCP) is NP-hard Bomze et al. (1999), finding all cliques is also NP-hard. Algorithms require an approximation or a shortcut due to enormous search space.
@@
Algorithm 1 describes the Loop and Clique Coarsening (LCC) process. Based on Tarjan Trajan (1974) and Clique Percolation (CP) Palla et al. (2005), LCC heuristically iterates the graph hierarchy using loop length and hierarchy depth as constraints for pruning. LCC finds cliques under depths less than $\sigma$ (lines 4-8) first, counting loops only when finding fewer or no cliques (lines 9-16). When updating graphs (lines 6 and 13), LCC reconstructs graphs to build the coarsening view. Mathematically, a new node $\bar{v}$ sums all node features in the structure (loop or clique) $S$ to represent it as follows:
$h_{\bar{v}}=\sum_{v_{i}\in S}h_{v_{i}},$ (2)
where $h_{*}$ denotes the feature of *.
Example 1.
We take Figure 2 as an example. Given an input graph $G_{i}$ (top left), we will get a reconstructed graph (top right) after coarsening loops and cliques.
As for the limitation of search complicity, we empirically set two constraints of loop length and hierarchy depth to prune.
Loop Length Constraint.
Not all loops are suitable for coarsening. Long loops may reveal two chains or sequences interacting (e.g., backbone in proteins Wang et al. (2023)). The semantics will change if coarsening happens. MSLgo sets a range $\delta$ for the maximum detection length (line 13). We default it to 6 to cover usual cases, such as squares and benzene rings.
Example 2.
The Loop Length Constraint in Figure 2 (bottom left) illustrates the case of $\delta=4,5$. Only loops with length $3<l\leq\delta$ are counted. It skips loops with length over $\delta$ and pushes triangles (3-clique) to Clique Coarsening.
Hierarchy Depth Constraint.
The NP-hard nature of finding all cliques determines the impossibility of exhaustive search, especially for large graphs. MSLgo takes advantage of the hierarchical characteristic of graphs, taking nodes with the highest degree (lines 3 and 7) hop by hop to find cliques. We set a distance $\sigma$ to control the depth of the recursion.
Example 3.
The Hierarchy Depth Constraint in Figure 2 (bottom center) shows choices of the yellow center node after graph coarsening with $\sigma=1$. It compresses the red clique and blue node in one hop while preserving the yellow one. Nodes at $d>\sigma$ stay still.
4.3 Line Graph Conversion
@@
This section tells why to build the Line Graph Conversion (LGC) view. Line graph is a usual method for edge representations in node-level tasks such as POI recommendation Chanpuriya et al. (2023). Edge augmentation is a considerable view to make a thorough representation integrating various architectures and scenarios Mo and Yamana (2023); Liu et al. (2023a). Therefore, we build LGC to attach an edge-central perspective in graph-level tasks. Firstly, the definition of line graphs is as follows.
Definition 3 (Line Graph).
@@
Given an undirected graph $G=\{\mathcal{V},\mathcal{E}\}$, $L(G)=\{L(\mathcal{E}),\mathcal{E}_{L(G)}\}$ is the line graph of $G$, where $L(\mathcal{E})$ and $\mathcal{E}_{L(G)}$ denote the node set and edge set of the line graph, and $L$ is a mapping function from the edge set of the original graph to the node set of the line graph, satisfying $\forall e_{ij}\in\mathcal{E},L(e_{ij})\in L(\mathcal{E})$, and $\forall e_{ij},e_{ik}\in\mathcal{E},(L(e_{ij}),L(e_{ik}))\in\mathcal{E}_{L(G)}(j\neq k)$.
@@
In plain words, a line graph turns the edges in the original graph into new vertices and the endpoints of two edges in the original graph into new edges. This explicit transformation can reduce the difficulty of relative position modeling, which is implicit in the node-central view of the original graph. We can build LGC through existing tools Hagberg et al. (2008) conveniently. For those graphs with no edge attributes, we sum features of endpoints for consistency following Equation 2. Moreover, we substitute dense graph inputs and large-scale graphs with multi-coarsening ones for lower time and space consumption.
@@
There exists relative position information loss once we treat graph structures node-like as in previous section. LGC is a supplementary of the coarsening view. We explain this from two aspects: different position information of neighborhood nodes and some hard-coarsening examples.
Different Position Information.
@@
From pre-experiments, we know that relative position suffers if we condense graph structures into one node. Though plain coarsening strategies decrease the graph scale, it makes positions vague, such as between Graphs A and B in Figure 1a. In essence, position information is the relative position among nodes connected by node-edge sequence. LGC models edge-central positional relationships, describing structural information by adding a new view to obtain a more comprehensive representation.
Example 4.
The Different Position Information in Figure 2 (bottom right) shows two similar graphs with a slight difference (the relative position of the edges is bold in red and blue). Coarsening the square will eliminate the difference, while LGC can enhance the diversity.
Hard Coarsening Examples.
MSLgo only focuses on loops and cliques. However, not all graphs contain either structures, such as long chains. A simple copy view contributes less if we leave these hard-coarsening examples alone. LGC transforms these graphs into a fittable form that is easy for postprocessing while preserving the original structural information.
Example 5.
The Hard Coarsening Examples in Figure 2 (bottom right) list two cases: claw-like structures (blue) and 4-node chains (red). Both are hard to coarsen since neither loop nor clique. After LGC, the former turns into a triangle, and the latter decreases the chain length. Both of the structures become easier for further processing.
4.4 Our Model: MSLgo
@@
After acquiring three views, we can build our model MSLgo based on U2GNN Nguyen et al. (2022). As shown in Figure 2, MSLgo takes $G_{i}$ as input and obtains the coarsening view $G_{co}$ and line graph conversion view $L(G_{i})$ with designed algorithms. For $G_{co}$, MSLgo calculates node embeddings $h$ according to Equation 1 and uses the summation readout function to obtain graph embedding $H_{co}$. Specifically, it is
$H_{co}=\sum_{v\in G_{co}}[h_{0,v}^{1};\cdots;h_{0,v}^{K}],$ (3)
@@
where $h_{0,v}^{k}$ denotes the initial embedding of node $v$ in layer $k$. For each layer $k$, MSLgo iterates $T$ steps aggregating $N$ sampling neighbors and passes $h_{T,v}^{k}$ to the next layer $k+1$.
@@
We use the same operation for the other two views, $G_{o}$ and $L(G_{i})$, and achieve corresponding embeddings, $H_{o}$ and $H_{lG}$. We apply concatenation across views to obtain the final embedding $H_{i}$ for the input graph $G_{i}$ as
$H_{i}=[H_{o};H_{co};H_{lG}].$ (4)
Apart from a single fully connected (FC) layer, we vary the classifier with a 2-layer MLP,
$\hat{y}_{i}=W_{2}(W_{1}H_{i}+b_{1})+b_{2},$ (5)
@@
where $W_{1}$ and $W_{2}$ are the weight matrices, and $b_{1}$ and $b_{2}$ are the bias parameters, adapting to the embedding dimension increase. The loss function is cross-entropy as follows:
$loss=-\sum_{i=1}^{M}y_{i}\log(\sigma(\hat{y}_{i})),$ (6)
@@
where $\sigma$ denotes the softmax function. For further implementation details, we suggest to refer to  Nguyen et al. (2022).
5 Experiments
5.1 Experimental Settings
Datasets.
We evaluate our approach on six widely used datasets from TUDataset Morris et al. (2020), including three social network datasets: COLLAB (CO), IMDB-BINARY (IB), and IMDB-MULTI (IM), and three bioinformatics datasets: D&D (DD), PTC_MR (PTC), and NCI1 (N1). Table 1 shows the details of the datasets.
Dataset # Graphs # Classes Avg.N Avg.E
CO 5000 3 74.49 2457.78
IB 1000 2 19.77 96.53
IM 1500 3 13.00 65.94
DD 1178 2 284.32 715.66
N1 4110 2 29.87 32.30
PTC 344 2 14.29 14.69

### v2

@@ 4.2.1. Algorithm Description
22:  return $G_{lcc}$
From pre-experiments, we know that treating graph structures node-like as a whole via graph coarsening is possible. Existing methods mainly pay attention to the hierarchical structure of graphs: some cluster nodes based on graph convolution to do graph coarsening (Ying et al., 2018), and some implement node condensing by iterative sampling (Lee et al., 2019). Aggregation of first-order adjacent nodes has intuitive interpretability, which means connections between neighborhoods, while a deeper coarsening shows a poor explanation. Moreover, in high-level coarsening, the main distinguishing node sets disappear into clusters, which leads to an ambiguity for topology. Thus, GRLsc achieves graph coarsening of loops and cliques with shallow coarse-grained clustering restricted by two hard constraints, which magnifies high-level structural information while preserving the characteristic nodes in graphs to the greatest extent.
We first give formal definitions of loops and cliques.
Definition 4.1 (Loop).
Given an undirected graph $G$, a path $G_{l}=\{v_{1}v_{2}\cdots v_{n}v_{n+1}\}$ is a loop if there is no repetitive node in $\{v_{1},\cdots,v_{n},v_{n+1}\}$ except $v_{1}=v_{n+1}$ and $(v_{k},v_{k+1})\in\mathcal{E}_{G}$ for $k=1,2,\cdots,n$. $G_{l}$ is a k-loop if $|\mathcal{V}_{G_{l}}|=k$.
Definition 4.2 (Clique).
@@
Coarsening requires counting all loops and cliques in the graph. However, according to the definitions above, loops are contained within cliques. For example, given a 4-clique marked $Cli=\{ABCD\}$, we can find four 3-loops: $\{ABC\}$, $\{ABD\}$, $\{ACD\}$, and $\{BCD\}$, and one 4-loop $\{ABCD\}$, except for one 4-clique, which is not we want, since identifying $Cli$ as a clique is much more straightforward than five loops. Moreover, since the Maximum Clique Problem (MCP) is NP-hard (Bomze et al., 1999), finding all cliques is also NP-hard. Algorithms require an approximation or a shortcut due to enormous search space.
@@
Algorithm 1 describes how the LCC Block works. Based on Tarjan (Trajan, 1974) and Clique Percolation (CP) (Palla et al., 2005), LCC heuristically iterates the graph hierarchy using loop length and hierarchy depth as constraints for pruning. LCC finds cliques under depths less than $\sigma$ (lines 3-13) first, counting loops only when finding fewer or no cliques (lines 14-21). When updating graphs, LCC reconstructs graphs to build the coarsening view. Mathematically, given an original graph $G_{o}=\{\mathcal{V},\mathcal{E}\}$, we aim to rebuild an intermediate coarsening graph $G_{lcc}$ with $n$ nodes where $n\ll|\mathcal{V}|$. In methods (Loukas and Vandergheynst, 2018; Loukas, 2019), supernode $v_{i}$ in $G_{lcc}$ aggregates nodes in $G_{o}$ according to node partitioning sets $\mathcal{P}=\{\mathcal{P}_{1},\cdots,\mathcal{P}_{n}\}$. We use indication matrix $M_{p}\in\{0,1\}^{n\times|\mathcal{V}|}$ to denote the aggregation based on partition set $\mathcal{P}$, where $m_{ij}=1$ if node $j$ in partition $\mathcal{P}_{i}$. We use a new adjacency matrix to represent $G_{lcc}$:
(2) $A_{lcc}=M_{p}AM_{p}^{T},$
where we multiply $M_{p}$ to the right to ensure the construction of the undirected graph. Here, we notice that $DEG=diag(A_{lcc})$ gives weight $a_{ii}$ for each supernode $i$, representing the sum of all node degrees corresponding to the partition $\mathcal{P}_{i}$. To directly extract and learn high-level structural information, we consider making a separation. We assign the weight of supernodes to 1 with $DEG^{-1}=diag(a_{ii}^{-1},\cdots,a_{nn}^{-1})$ instead of summation. The normalized adjacency matrix is
(3) $\bar{A}_{lcc}=diag(a_{ii}^{-1},\cdots,a_{nn}^{-1})(M_{p}AM_{p}^{T}),$
leaving each supernode equal contribution to the high-level structure initially. Similarly, we can define the node features of $G_{lcc}$ using $M_{p}$:
(4) $H_{lcc}=M_{p}H,$
where $H_{*}$ indicates the feature representation of *. This equation is equivalent to summing the nodes according to the partition sets.
So far, the focus of graph coarsening falls on how to obtain the partitioning set $\mathcal{P}$. LCC finds partitioning sets by searching for loops and cliques, where each $\mathcal{P}_{i}$ represents a loop or a clique. In other words, the process of LCC is to find a linear transformation matrix $A_{L}$ about $A$, such that
(5) $A^{\prime}=A_{L}AA_{L}^{T}=\begin{bmatrix}\mathcal{P}_{1}&E_{12}&\cdots&E_{1n}\\ E_{21}&\mathcal{P}_{2}&\ddots&\vdots\\ \vdots&\ddots&\ddots&\vdots\\ E_{n1}&\cdots&\cdots&\mathcal{P}_{n}\end{bmatrix},$
where the diagonal $\mathcal{P}_{i}=\{1\}^{k_{i}\times k_{i}}$ is the partitioning set containing $k_{i}$ nodes, and the non-diagonal $E_{ij}\in\{0,1\}^{k_{i}\times k_{j}}$ records the connection between partition $\mathcal{P}_{i}$ and partition $\mathcal{P}_{j}$. Next, by mapping each submatrix to 0 or 1 with the indicator function $\mathcal{I}(X)$, we can get the same result as Equation 2, i.e,
(6) $A_{lcc}=\mathcal{I}(A^{\prime}),$
(7) $\mathcal{I}(X)=\begin{cases}k_{i}&X=\mathcal{P}_{i}\\ 1&\exists x\in X,X=E_{ij},x=1\\ 0&\text{otherwise}\end{cases}.$
Figure 3. LCC Block contains two procedures: clique coarsening and loop coarsening. We design two constraints, hierarchy depth, and loop length, to control the graph coarsening process. A-D in the blue enclosure shows the details of the clique coarsening under hierarchy depth constraint, ①-③ in the red the loop coarsening with loop length constraint.
Figure 3 further depicts the workflow of the LCC Block. We design two coarsening algorithms with hierarchy depth constraint for clique and loop length constraint for loop. We go into details one by one.
4.2.2. Clique Coarsening with Hierarchy Depth Constraint.
The NP-hard nature of finding all cliques determines the impossibility of exhaustive search, especially for large graphs. GRLsc takes advantage of the hierarchical characteristic of graphs, taking nodes with the highest degree (Alg 1 line 2) hop by hop to find cliques. We set a distance $\sigma$ to control the depth of the recursion. A-D inside the blue closure in Figure 3 demonstrates the clique coarsening process with hierarchy depth constraints. (A) Given an example input graph with a clique structure, consider the node $M$ with the highest degree as the center node. (B) Search for possible cliques formed by connections within 1-hop neighbors of the central node and coarsen the found clique structure to rebuild a new node while preserving the central node. In step C, the central node $M$ remains unchanged, and the supernode $A=\{A_{1},A_{2},A_{3},M\}$. (C) Set the hierarchy depth constraint $\sigma$ and stop the coarsening of the current central node when the search range exceeds the limitation. If $\sigma=1$, the searching process simplifies to DFS. (D) If there exist nodes having 1-hop neighbors not searched yet (Alg 1 line 7), switch the central node and repeat the flow A-D until all nodes and edges are covered. For example, $C_{1},C_{2}$ are around node $B$, located outside the $\sigma$-hop of the central node $M$. After switching the central node to $B$, we can continue to search for the cliques.
4.2.3. Loop Coarsening with Loop Length Constraint.
Not all loops are suitable for coarsening. Long loops may reveal two chains or sequences interacting (e.g., backbone in proteins (Wang et al., 2023)). The semantics will change if coarsening happens. GRLsc sets a range $\delta$ for the maximum detection length (Alg  1 line 17). We default it to 6 to cover usual cases, such as squares and benzene rings. ①-③ inside the red closure in Figure 3 gives the loop coarsening process with loop length constraint. ① Given an example graph with loops of different lengths. ② Pick the starting node and transform the graph into a DFS sequence. ③ Set the loop length constraint $\delta$. When the sequence length exceeds the constraint, we can make the pruning on that chain. For example, the sample graph contains two 4-loop: $\{ABEF\}$ and $\{BCDE\}$ and one 6-loop: $\{ABCDEF\}$. When $\delta=4,5$, only $\{ABEF\}$ and $\{BCDE\}$ will be selected, forming a coarsening graph consisting of two supernodes. When $\delta=6$, $\{ABCDEF\}$ is added to the candidate set, replacing two loops above to construct one supernode. Although the triangle does not appear in the sample graph, it is a plain structure in the real world. We can handle it either as a 3-loop or a 3-clique.
4.2.4. Time Complexity
According to the previous description, LCC Block contains two parts, clique and loop coarsening, and the total time complexity of the algorithm is the sum of them. Given the input graph $G=\{\mathcal{V},\mathcal{E}\}$, we search for cliques following the hierarchy depth $\sigma$, marking nodes visited if no possible neighbors can form the new cliques. The time complexity of clique coarsening is $\mathcal{O}(\mathcal{V}+\mathcal{E})$. $\sigma$ only gives an order for searching and does not influence the linear computational cost. As for loop coarsening, we traverse along the DFS sequence and only keep loops under length constraints $\delta$. The time complexity of loop coarsening is $\mathcal{O}(\delta\mathcal{V}+\mathcal{E})$. For each node, we look back within $\delta$-level seeking loops satisfying the needs. Since $\delta$ is relatively small, the computation stays linear, $\mathcal{O}(\mathcal{V}+\mathcal{E})$. Therefore, the time complexity of the final LCC Block is $\mathcal{O}(2(\mathcal{V}+\mathcal{E}))$, achieving a linear growth in the scale of the input graph.
4.2.5. Limitation
Both loops and cliques are common structures in real-world graph networks. Understanding these structures is beneficial to thoroughly learning the network. LCC Block focuses on solving loops and cliques, so it is limited to other networks where these structures do not exist (e.g., long chains). To effectively alleviate this limitation of LCC, we introduce LGC Block next section 4.3 to further supplement the structural information through the conversion.
4.3. Line Graph Conversion
Figure 4. LGC Block workflow. Left side A-B shows the line graph conversion with hard coarsening examples (a.1 and a.2). Right side ①-③ illustrates how the input graphs with different position information are distinguished by LGC but not LCC.
@@ 4.3. Line Graph Conversion
This section tells why to build the LGC view. Line graph is a usual method for edge representations in node-level tasks such as POI recommendation (Chanpuriya et al., 2023). Edge augmentation is a considerable view to make a thorough representation integrating various architectures and scenarios (Mo and Yamana, 2023; Liu et al., 2023a). Therefore, we build LGC to attach an edge-central perspective in graph-level tasks. Firstly, the definition of line graphs is as follows.
Definition 4.3 (Line Graph).
@@
Given an undirected graph $G_{o}=\{\mathcal{V},\mathcal{E}\}$, $L(G)=\{L(\mathcal{E}),\mathcal{E}_{L(G)}\}$ is the line graph of $G$, where $L(\mathcal{E})$ and $\mathcal{E}_{L(G)}$ denote the node set and edge set of the line graph, and $L$ is a mapping function from the edge set of the original graph to the node set of the line graph, satisfying $\forall e_{ij}\in\mathcal{E},L(e_{ij})\in L(\mathcal{E})$, and $\forall e_{ij},e_{ik}\in\mathcal{E},(L(e_{ij}),L(e_{ik}))\in\mathcal{E}_{L(G)}(j\neq k)$.
@@
In plain words, a line graph turns the edges in the original graph into new vertices and the endpoints of two edges in the original graph into new edges. This explicit transformation can reduce the difficulty of relative position modeling, which is implicit in the node-central view of the original graph. We can build LGC through existing tools (Hagberg et al., 2008) conveniently. For those graphs with no edge attributes, we sum features of endpoints for consistency following Equation 4. In other words, we takes edges as the partitioning set and further aggregates features to build the LGC view. Mathematically, given an input graph $G_{o}$ and its feature vector $H$, LGC computes:
(8) $H_{lgc}=M_{e}[L(H_{lcc})]=M_{e}[L(M_{p}H)],$
where $L(\cdot)$ is the conversion function according to definition 4.3, $M_{e}$ is the indication matrix with edge partitions, and $H_{lcc}$ is the feature representations output by Equation 4.
@@
In previous sections, we know LCC is limited if no loops and cliques exist, and the model suffers relative position information loss if we treat structures node-like. LGC is a supplementary of the coarsening view. We explain this from two aspects: some hard-coarsening examples and different position information of neighborhood nodes. Figure 4 shows the details.
4.3.1. LGC Block with Hard Coarsening Examples.
GRLsc only focuses on loops and cliques. However, not all graphs contain either structures, such as long chains. A simple copy view contributes less if we leave these hard-coarsening examples alone. LGC transforms these graphs into a fittable form that is easy for postprocessing while preserving the original structural information. For example, A-B in Figure 4 shows the LGC workflow with some hard coarsening examples. (A) Given a sample input graph without loops and cliques, we exclude two hard-coarsening structures: (a.1) Claw-like structure, consisting of a central node and three (or more) independent nodes connected to it. (a.2) Long chain with $n(=3)$ nodes connected one by one. Neither of them can be handled efficiently by LCC. (B) Convert structures through LGC Block to build an edge-centric view. In a claw-like structure, all edges are connected through the central node, forming a new clique after conversion. The long chain with $n$ nodes can decline the length to $n-1$ in one LGC calculation, reducing the scale effectively. GRLsc applies one LGC step on the intermediate coarsening graph of LCC, retaining the newly generated structures after conversion.
4.3.2. Different Position Information.
@@
From pre-experiments, we know that relative position suffers if we condense graph structures into one node. Though plain coarsening strategies decrease the graph scale, it makes positions vague, such as between Graphs A and B in Figure 1. In essence, position information is the relative position among nodes connected by node-edge sequence. LGC models edge-central positional relationships, describing structural information by adding a new view to obtain a more comprehensive representation. We use ①-③ on the right of Figure 4 for illustration. ① Given two sample input graphs, each contains six nodes, four of which form a square, and the other two (denoted by A and B in the figure) connect to the square. The two sample graphs differ only in connection positions, one at adjacent nodes in the square and the other at diagonal nodes. ② If the square is identified as a 4-loop and coarsened to a supernode by LCC, it can not distinguish the two graphs by connection cases. ③ Realize the conversion to edge-central view by LGC, which can preserve different position information and differentiate graphs with subtle structural differences. As shown in Figure 4, LGC can identify the local claw-like structure containing node $B$ and convert it into a new triangle. We will get two LGC views with different positions of the triangle of node $B$ (in blue closure) adjacent or opposite to the triangle of $A$.
4.3.3. Limitation
Although LGC is complementary to LCC, it still has limitations. LGC can help LCC deal with graphs that don’t have loops and cliques, but line graph conversion is mostly a complication of the graph (increasing the scale). For a fully connected graph $G$ with $n$ nodes, the line graph $L(G)$ contains $\frac{n(n-1)}{2}$ nodes. It is an unacceptable growth rate at a quadratic level. LGC has poor adaptability to dense graphs, and its performance decreases when the number of edges in the intermediate coarsening graph by LCC far exceeds the number of nodes.
4.4. Our Model: GRLsc
@@
After acquiring three views, we can build our model GRLsc based on U2GNN (Nguyen et al., 2022). As shown in Figure 2, GRLsc takes $G_{i}$ as input and obtains the coarsening view $G_{lcc}$ and line graph conversion view $G_{lgc}$ with LCC and LGC Block, respectively. For $G_{lcc}=\{\mathcal{V}_{lcc},\mathcal{E}_{lcc}\}$, GRLsc calculates node embeddings $h$ according to Equation 1 and Equation 4 and uses the summation readout function to obtain graph embedding $H_{lcc}$. Specifically, it is
(9) $H_{lcc}=\sum_{v\in\mathcal{V}_{lcc}}[h_{0,v}^{1};\cdots;h_{0,v}^{K}],$
@@
where $h_{0,v}^{k}$ denotes the initial embedding of node $v$ in layer $k$. For each layer $k$, GRLsc iterates $T$ steps aggregating $N$ sampling neighbors and passes $h_{T,v}^{k}$ to the next layer $k+1$.
@@
We use the same operation for the other two views, $G_{o}$ and $G_{lgc}$, and achieve corresponding embeddings, $H_{o}$ and $H_{lgc}$. We apply concatenation across views to obtain the final embedding $H_{i}$ for the input graph $G_{i}$ as
(10) $H_{i}=[H_{o};H_{lcc};H_{lgc}].$
After that, we feed the embedding $H_{i}$ to a single fully connected (FC) layer.
(11) $\hat{y}_{i}=WH_{i}+b,$
@@
where $W$ is the weight matrix, and $b$ is the bias parameter, adapting to the embedding dimension increase. The loss function is cross-entropy as follows:
(12) $loss=-\sum_{i=1}^{M}y_{i}\log(\sigma(\hat{y}_{i})),$
@@
where $\sigma$ denotes the softmax function. For further implementation details, we suggest to refer to  (Nguyen et al., 2022).
5. Experiments
5.1. Experimental Settings
5.1.1. Datasets.